# BRAiSE: synthetic polarisation in RMHD AGN jet simulations

### v1

@@ Abstract
doi:
10.1017/pasa.2020.32keywords:
@@ 1 Introduction
Meenakshi et al. (2023) have published the most advanced simulation of polarised radio emission in AGN jets and lobes to date. This method uses a hybrid grid and particle method by Vaidya et al. (2018) to track the evolution of each particle as it is advected by the fluid on the simulation grid. The energy distribution of the particles is followed in-situ by solving the relativistic cosmic ray transport equation. The surface brightness is then ray-traced in post-processing. The authors study radio lobes of intermediate sizes of up to approximately kpc in total length. This continues a tradition of in-situ treatment of electrons in AGN jet simulations (e.g. Jones et al., 1999; Tregillis et al., 2004; Jones & Kang, 2005; Nolting et al., 2022).
Kramer et al. (2024) use a similar method in their RMHD simulations of parsec-scale AGN jets. The authors use Lagrangian particles in their simulations and follow the energy distribution of the particles in a similar manner to Vaidya et al. (2018), however, they post-process their simulated data using the RADMC-3D code (MacDonald & Nishikawa, 2021). The RADMC-3D code calculates the full radiative transfer of the Stokes’ parameters as based on the methods of Jones & O’Dell (1977a, b). We do not consider the full radiative transfer of the Stokes’ parameters in this work.
In this paper we introduce a new method of calculating polarised radio emission, building on the PRAiSE (Particles + Radio AGN in Semi-analytic Environments) framework (Yates-Jones et al., 2022; Turner et al., 2018a). The PRAiSE method is a hybrid approach, using Lagrangian particles that are advected by the fluid in a grid-based simulation. These particles record the history of the local fluid values (pressure, density, velocity, magnetic field) encountered along their trajectory and the time since the particle last passed through a shock (using three thresholds, see Yates-Jones et al., 2022, for more details). The particle energy distributions and the associated synchrotron emission are calculated in post-processing, accounting for adiabatic and IC losses. The evolution of particle energy is found by interpolating the emitting particle Lorentz factor backwards in time to the last acceleration event experienced by that particle. In contrast to the methods used by Vaidya et al. (2018) and Kramer et al. (2024), our new method allows for more flexibility in our parameter choice (e.g. injection index, redshift) as the computationally efficient calculation is performed in post-processing, rather than in-situ.
@@ 2.1 Synchrotron emission
where is the number density of radiating particles, is the mass of the radiating particles, is the number density of thermal particles, and is the mass of the thermal particles. The volume filling factor then is:
@@ 3.1 Simulations
The simulations in this paper are carried out using the PLUTO astrophysical fluid dynamics code (version 4.3; Mignone et al., 2007) using the relativistic magnetohydrodynamics (RMHD) physics and relativistic hydrodynamics (RHD) physics modules. The HLLD Riemann solver is used, as well as 2nd order Runge-Kutta timestepping and linear reconstruction. The condition is controlled by Powell’s eight-wave formulation (Powell, 1997; Powell et al., 1999), which is used to prevent nonphysical changes in the magnetic field structure. Radiative cooling is not considered in these simulations; all radiative losses are calculated in post-processing as they do not significantly affect jet dynamics in the evolutionary stages considered in this work (Hardcastle, 2018).
These simulations were carried out on a three-dimensional Cartesian grid centered at . Each dimension contains five grid patches; a uniform grid with grid cells from kpc to ensure jet injection is sufficiently resolved; two stretched grid patches from kpc; and two stretched grid patches from kpc. The stretched grid patches contain and cells respectively. All simulation grid boundaries are periodic.
@@ 3.2 Dynamics
Fig. 1 shows this difference in the morphology of our three simulated jets, plotted at the same total lobe length. The lobes in simulations RAG-B0 and RAG-B16 are almost identical; since the jet magnetic energy is quite low ( percent of the total jet power), the magnetic field does not change the dynamics of the jets in simulation RAG-B16 significantly (e.g. Hardcastle & Krause, 2014). The lobe in simulation RAG-B327 is narrower by a factor of and reaches the total length of kpc times faster than simulations RAG-B0 and RAG-B16 and times faster than simulation RAC-B327. Since simulation RAC-B327 has a more dense environment, the expansion of the lobe along the jet axis is inhibited, resulting in a narrow jet head and wide equatorial region of the lobe.
Our simulations are comparable to the moderate-powered jets in Mukherjee et al. (2020)’s simulations D (RAG-B16) and F (RAG-B327 and RAC-B327), but shown here on larger scales as they have been evolved for longer. The difference in lobe width and jet stability is the same on these larger scales as that reported by Mukherjee et al. (2020) on galactic scales; the jet in RAG-B16 is disrupted at Myr by Kelvin-Helmholtz instabilities, whereas in simulations RAG-B327 and RAC-B327 the jets remain relatively stable to these instabilities due to their higher magnetisation. These simulations confirm the result of Mukherjee et al. (2020) that higher powered, faster jets, with strong magnetic fields are the most stable RMHD jets.
@@ 3.3 Surface brightness
We calculate the surface brightness images in the plane of the sky following the method outlined in (Yates-Jones et al., 2022). The images have been convolved with a 2D Gaussian beam with FWHM of at a redshift of and at (this corresponds to the same physical beam size of kpc). In Fig. 6, we show the surface brightness images for simulations RAG-B0 and RAG-B16 at Myr; for the latter RMHD simulation, we include both PRAiSE and BRAiSE images and the difference between them. In Fig. 7 we plot the surface brightness images and their difference map for simulation RAG-B327 at Myr, and similarly plot these images at Myr for simulation RAC-B327 in Fig. 8.
@@ 3.4 Radio morphology
To ensure that these results are not an artefact of calculating the surface brightness using the Lagrangian particles, we have tested the convergence of the calculated emissivity for different numbers of particles injected onto the simulation grid over time. We find that our particle injection rate of particles per Myr is more than sufficient to recover the total lobe emissivity, which is constant for an injection rate above particles per Myr. The flux density of the fast jet slowly decreases as more particles are added, however, this contributes a small fraction of the total emission. We note that local brightness inhomogeneities in the surface brightness images are smoothed out somewhat as the number of particles increase, and we caution against over-interpreting detailed structure in the surface brightness.
@@ 4.1 Fractional polarisation without Faraday rotation
The differences in jet dynamics due to jet magnetisation impact the polarisation properties of our simulated sources. In Fig. 11 we plot the fractional polarisation overlaid with contours corresponding to the Stokes I emission (from Figs. 6, 7, 8) for each of our RMHD simulations. The length of the vectors corresponds to the fractional polarisation and the direction indicates the inferred ‘magnetic field direction’, , following Hardcastle & Krause (2014). Two frequencies are plotted ( GHz and GHz), at low resolution (beam FWHM of kpc, corresponding to at and at ) and high resolution (beam FWHM of kpc, corresponding to at and at ) in the left and right hand columns for each frequency respectively. For each frequency and resolution, we plot low () and high () redshift images in the first four and latter four columns respectively. We have held the physical pixel and beam sizes constant across redshifts to compare the effect of IC losses on the fractional polarisation. In the upper three rows, we plot the raw fractional polarisation images, and in the lower three rows, we plot the depolarised by Faraday rotation fractional polarisation images (discussed in Section 4.2 below).
@@
In Fig. 12, we plot the integrated fractional polarisation, ; i.e. the fractional polarisation that would be measured if the source was unresolved. We include three different viewing angles to the source; to the plane of the sky. In this section, we do not consider the effects of Faraday rotation (these effects are discussed in Section 4.2).
@@ 4.2 Depolarisation due to Faraday rotation
Hardcastle & Krause (2014) study the Laing-Garringon effect in their simulations using a depolarisation ratio (DPR), measured as a ratio of the depolarisation of the jet and counterjet. This depolarisation for each jet is the ratio of polarised flux density (with Faraday rotation applied) for a low and high frequency respectively; Hardcastle & Krause (2014) choose frequencies of and GHz. The DPR should be greater than 1 if the jet (pointed towards the observer) is less depolarised than its counterjet (pointed away from the observer), as expected from the Laing-Garrington effect (Laing, 1988; Garrington et al., 1988). This is due to the fact that the jet that is pointed towards the observer will have lower Faraday rotation measure values as there will be a shorter path length through the shocked shell of ambient medium that surrounds the jet cocoon (e.g. Jerrim et al., 2024).
@@ 5 Conclusions
High resolution surface brightness images produced using BRAiSE show clumps due to the complex magnetic field structure in the lobe. In general, BRAiSE produces sources with greater total luminosity due to the variation of magnetic field strength across the lobes. We simulate three sources; one that resembles a lobed FR-I source, and two that resemble FR-II sources. BRAiSE produces a diffuse hotspot region in our FR-II-like sources, since the magnetic energy at the jet head is spread out due to kink instabilities in the jet. BRAiSE also produces strong emission in a flaring point for these sources, which affects their classification using the FR index. We find that for PRAiSE, the FR index classifies each of our sources as an FR-II. However, for BRAiSE, these sources are more borderline. For our highly magnetised jets, BRAiSE predicts significantly lower FR indices than PRAiSE. We conclude that magnetic field information is required to interpret the FR morphology, and that our simulations show unique shock structures with the presence of both a flaring point and a hotspot that result in a bimodal surface brightness distribution. This unique shock structure indicates that we have not simulated true FR-II sources, and that our jet setup is missing the full set of physics required to stablise the jet on large scales (e.g. jet rotation; Bodo et al., 2016, 2019).
@@ 6 Acknowledgements
This research was carried out using the high-performance computing clusters provided by Digital Research Services, IT Services at the University of Tasmania. This work has made use of data from The Three Hundred collaboration (https://www.the300-project.org) which benefits from financial support of the European Union’s Horizon 2020 Research and Innovation programme under the Marie Skłodowskaw-Curie grant agreement number 734374, that is the LACEGAL project. The Three Hundred simulations used in this paper have been performed in the MareNostrum Supercomputer at the Barcelona Supercomputing Center, thanks to CPU time granted by the Red Española de Supercomputación. LJ thanks the University of Tasmania for an Australian Government Research Training Program (RTP) Scholarship. SS, CP and PYJ acknowledge funding by the Australian Research Council via grant DP240102970. We thank Knapen et al. (2022) for their guide to writing astronomy papers. We acknowledge the support of the developers providing the Python packages used in this paper: Astropy (Astropy Collaboration et al., 2022), JupyterLab (Kluyver et al., 2016), Matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), and SciPy (Virtanen et al., 2020).

### v2

@@ Abstract
keywords:
magnetohydrodynamics – galaxies: active – galaxies: jets – radio continuum: galaxiesdoi:
@@ 1 Introduction
Meenakshi et al. (2023) have published the most advanced RMHD simulations of polarised radio emission in AGN jets and lobes to date. This method uses a hybrid grid and particle method by Vaidya et al. (2018) to track the evolution of each particle as it is advected by the fluid on the simulation grid. The energy distribution of the particles is followed in-situ by solving the relativistic cosmic ray transport equation. The synchrotron emissivity and full Stokes parameters are calculated within the simulations, and the surface brightness images are then ray-traced in post-processing. The authors study radio lobes of intermediate sizes of up to approximately kpc in total length. This continues a tradition of in-situ treatment of electrons in AGN jet simulations (e.g. Jones et al., 1999; Tregillis et al., 2004; Jones & Kang, 2005; Nolting et al., 2022).
Kramer et al. (2024) use a similar method in their RMHD simulations of parsec-scale AGN jets. The authors use Lagrangian particles in their simulations and follow the energy distribution of the particles in a similar manner to Vaidya et al. (2018), however, they post-process their simulated data (including calculating the polarised emission) using the RADMC-3D code (MacDonald & Nishikawa, 2021). The RADMC-3D code calculates the full radiative transfer of the Stokes’ parameters as based on the methods of Jones & O’Dell (1977a, b). We do not consider the full radiative transfer of the Stokes’ parameters in this work. This hybrid method incorporates a mixture of in-situ and ex-situ treatments (in-situ particles/energy distribution, ex-situ emissivity), in contrast to a fully post-processed method (e.g. Fromm et al., 2016) or fully in-situ method (e.g. Meenakshi et al., 2023).
In this paper we introduce a new method of calculating polarised radio emission, building on the PRAiSE (Particles + Radio AGN in Semi-analytic Environments) framework (Yates-Jones et al., 2022; Turner et al., 2018a). The PRAiSE method is also a hybrid approach, using Lagrangian particles that are advected by the fluid in a grid-based simulation to evolve the particle energy distribution and calculate synchrotron emissivity ex-situ. These particles record the history of the local fluid values (pressure, density, velocity, magnetic field) encountered along their trajectory and the time since the particle last passed through a shock (using three thresholds, see Yates-Jones et al., 2022, for more details). The particle energy distributions and the associated synchrotron emission are calculated in post-processing, rather than within the simulation, and account for (re-)acceleration due to strong shocks, and subsequent adiabatic, synchrotron, and IC losses. The evolution of particle energy is found by interpolating the emitting particle Lorentz factor backwards in time to the last acceleration event experienced by that particle. In contrast to the methods used by Vaidya et al. (2018) and Kramer et al. (2024), our method allows for more flexibility in our parameter choice (e.g. injection index, redshift) as the computationally efficient particle energy distribution calculation is performed in post-processing, rather than in-situ.
@@ 2.1 Synchrotron emission
where is the number density of radiating particles, is the mass of lobe plasma associated with each radiating particle (i.e., the sum of electron mass and positively charged particle mass), is the number density of thermal particles, and is the mass of the thermal particles. The volume filling factor then is:
@@ 3.1 Simulations
The simulations in this paper are carried out using the PLUTO astrophysical fluid dynamics code (version 4.3; Mignone et al., 2007) using the relativistic magnetohydrodynamics (RMHD) physics and relativistic hydrodynamics (RHD) physics modules. The HLLD Riemann solver is used, as well as 2nd order Runge-Kutta timestepping with a Courant-Friedrichs-Lewy (CFL) number of and linear reconstruction. The condition is controlled by Powell’s eight-wave formulation (Powell, 1997; Powell et al., 1999), which is used to prevent nonphysical changes in the magnetic field structure. Radiative cooling is not considered in these simulations; all radiative losses are calculated in post-processing as they do not significantly affect jet dynamics in the evolutionary stages considered in this work (Hardcastle, 2018).
These simulations were carried out on a three-dimensional Cartesian grid centred at . Each dimension contains five grid patches; a central uniform grid with grid cells from kpc with a resolution of kpc/cell to ensure jet injection is sufficiently resolved; two stretched grid patches from kpc; and two stretched grid patches from kpc. The stretched grid patches contain and cells respectively, with typical resolutions of kpc/cell at kpc and kpc/cell at kpc. The total grid size is with periodic boundary conditions at each outer boundary. The unit values of length, velocity and density in our simulations are kpc, , and g/cm3 respectively.
@@ 3.2 Dynamics
Fig. 1 shows this difference in the morphology of our three simulated jets, plotted at the same total lobe length. The lobes in simulations RAG-B0 and RAG-B16 are almost identical; since the jet magnetic energy is quite low ( percent of the total jet power), the magnetic field does not change the dynamics of the jets in simulation RAG-B16 significantly (e.g. Hardcastle & Krause, 2014). The lobe in simulation RAG-B327 is narrower by a factor of and reaches the total length of kpc times faster than simulations RAG-B0 and RAG-B16 and times faster than simulation RAC-B327. Since simulation RAC-B327 has a denser environment, the expansion of the lobe along the jet axis is inhibited, resulting in a narrow jet head and wide equatorial region of the lobe.
Our simulations are comparable to the moderate-powered jets in Mukherjee et al. (2020)’s simulations D (RAG-B16) and F (RAG-B327 and RAC-B327) but shown here on larger scales as they have been evolved for longer. The difference in lobe width and jet stability is the same on these larger scales as that reported by Mukherjee et al. (2020) on galactic scales; the jet in RAG-B16 is disrupted at Myr by Kelvin-Helmholtz instabilities, whereas in simulations RAG-B327 and RAC-B327 the jets remain relatively stable to these instabilities due to their higher magnetisation. These simulations confirm the result of Mukherjee et al. (2020) that higher powered, faster jets, with strong magnetic fields are the most stable RMHD jets.
@@ 3.3 Surface brightness
We calculate the surface brightness images in the plane of the sky following the method outlined in (Yates-Jones et al., 2022). The images have been convolved with a 2D Gaussian beam with FWHM of at a redshift of and at (this corresponds to the same physical beam size of kpc). The choice of this beam FWHM is representative of high-resolution SKA pathfinder instruments (Morabito et al., 2022). In Fig. 6, we show the surface brightness images for simulations RAG-B0 and RAG-B16 at Myr; for the latter RMHD simulation, we include both PRAiSE and BRAiSE images and the difference between them. In Fig. 7 we plot the surface brightness images and their difference map for simulation RAG-B327 at Myr, and similarly plot these images at Myr for simulation RAC-B327 in Fig. 8.
@@ 3.4 Radio morphology
To ensure that these results are not an artefact of calculating the surface brightness using the Lagrangian particles, we have assessed the convergence of the calculated emissivity for different numbers of particles injected onto the simulation grid over time. We find that our particle injection rate of particles per Myr is more than sufficient to recover the total lobe emissivity, which is constant for an injection rate above particles per Myr. The flux density of the fast jet slowly decreases as more particles are added, however, this contributes a small fraction of the total emission. We note that local brightness inhomogeneities in the surface brightness images are smoothed out somewhat as the number of particles increase, and we caution against over-interpreting detailed structure in the surface brightness.
@@ 4.1 Fractional polarisation without Faraday rotation
The differences in jet dynamics due to jet magnetisation impact the polarisation properties of our simulated sources. In Fig. 11 we plot the fractional polarisation overlaid with contours corresponding to the Stokes I emission (from Figs. 6, 7, 8) for each of our RMHD simulations. The length of the vectors corresponds to the fractional polarisation and the direction indicates the inferred ‘magnetic field direction’, , following Hardcastle & Krause (2014). Two frequencies are plotted ( GHz and GHz), at low resolution (beam FWHM of kpc, corresponding to at and at ) and high resolution (beam FWHM of kpc, corresponding to at and at ) in the left- and right-hand columns for each frequency respectively. The low resolution beam FWHM is consistent with beam FWHMs in current and upcoming polarisation surveys, which typically range between (see Table 3 of Jerrim et al., 2024). For each frequency and resolution, we plot low () and high () redshift images in the first four and latter four columns respectively. We have held the physical pixel and beam sizes constant across redshifts to compare the effect of IC losses on the fractional polarisation. In the upper three rows, we plot the raw fractional polarisation images, and in the lower three rows, we plot the depolarised by Faraday rotation fractional polarisation images (discussed in Section 4.2 below).
@@
In Fig. 12, we plot the integrated fractional polarisation, ; i.e. the fractional polarisation that would be measured if the source was unresolved. We include three different viewing angles to the source: and to the plane of the sky. In this section, we do not consider the effects of Faraday rotation (these effects are discussed in Section 4.2).
@@ 4.2 Depolarisation due to Faraday rotation
Hardcastle & Krause (2014) study the Laing-Garrington effect in their simulations using a depolarisation ratio (DPR), measured as a ratio of the depolarisation of the jet and counterjet. This depolarisation for each jet is the ratio of polarised flux density (with Faraday rotation applied) for a low and high frequency respectively; Hardcastle & Krause (2014) choose frequencies of and GHz. The DPR should be greater than 1 if the jet (pointed towards the observer) is less depolarised than its counterjet (pointed away from the observer), as expected from the Laing-Garrington effect (Laing, 1988; Garrington et al., 1988). This is due to the fact that the jet that is pointed towards the observer will have lower Faraday rotation measure values as there will be a shorter path length through the shocked shell of ambient medium that surrounds the jet cocoon (e.g. Jerrim et al., 2024).
@@ 5 Conclusions
High resolution surface brightness images produced using BRAiSE show clumps due to the complex magnetic field structure in the lobe. In general, BRAiSE produces sources with greater total luminosity due to the variation of magnetic field strength across the lobes. We simulate three sources; one that resembles a lobed FR-I source, and two that resemble FR-II sources. BRAiSE produces a diffuse hotspot region in our FR-II-like sources, since the magnetic energy at the jet head is spread out due to kink instabilities in the jet. BRAiSE also produces strong emission in a flaring point for these sources, which affects their classification using the FR index. We find that for PRAiSE, the FR index classifies each of our sources as an FR-II. However, for BRAiSE, these sources are more borderline. For our highly magnetised jets, BRAiSE predicts significantly lower FR indices than PRAiSE. We conclude that magnetic field information is required to interpret the FR morphology, and that our simulations show unique shock structures with the presence of both a flaring point and a hotspot that result in a bimodal surface brightness distribution. This unique shock structure indicates that we have not simulated true FR-II sources, and that our jet setup is missing the full set of physics required to stabilise the jet on large scales (e.g. jet rotation; Bodo et al., 2016, 2019).
@@ 6 Acknowledgements
LJ thanks the University of Tasmania for an Australian Government Research Training Program (RTP) Scholarship. SS, CP, and PYJ acknowledge funding by the Australian Research Council via grant DP240102970. This research was carried out using the high-performance computing clusters provided by Digital Research Services, IT Services at the University of Tasmania. This work has made use of data from The Three Hundred collaboration (https://www.the300-project.org) which benefits from financial support of the European Union’s Horizon 2020 Research and Innovation programme under the Marie Skłodowskaw-Curie grant agreement number 734374, that is the LACEGAL project. The Three Hundred simulations used in this paper have been performed in the MareNostrum Supercomputer at the Barcelona Supercomputing Center, thanks to CPU time granted by the Red Española de Supercomputación. We thank Knapen et al. (2022) for their guide to writing astronomy papers. We acknowledge the support of the developers providing the Python packages used in this paper: Astropy (Astropy Collaboration et al., 2022), JupyterLab (Kluyver et al., 2016), Matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), and SciPy (Virtanen et al., 2020).